OpenStudy (anonymous):

HELP ME OUT WITH MATH MEDALS WILL BE GIVEN OUT 1. y minus five over three= 1 (1 point) –2 8 18 6 2. 6x + 29 = 5 (1 point) –4 –18 204 –144 3. Combine like terms: –21a + 16a (1 point) –5a 5a 37a –37a Simplify the expression. 4. 9a – b – 2a – 10b (1 point) –7a + 11b 11a + 9b –11a – 9b 7a – 11b 5. 5(x + 10) + x (1 point) 6x + 15 5x + 15 6x + 50 4x – 50

OpenStudy (skullpatrol):

$$\Huge\text{ONE question at a time please.}$$

OpenStudy (anonymous):

1. y minus five over three= 1 (1 point) –2 8 18 6

OpenStudy (skullpatrol):

$$\Huge\text{ONE question per post please}$$

OpenStudy (anonymous):

2, 6x+29=5 6x=5-29 6x=-24 x=-24:6 x=-4